Can you please help me simplify -5x^0 please? Thanks!
Found 2 solutions by tutorcecilia, fractalier:
Answer by tutorcecilia(2144)

Any number raised to the zero power is 1.
So -5x^0 = (-5)(1) = -5

Answer by fractalier(2101)

Since anything to the zero power is 1, we have
-5x^0 = -5(1) = -5